OpenStudy (bekkah323):

find the function values g(x)=-2x+5; g(0) and g(-1)

OpenStudy (whpalmer4):

just write out the function, replace x wherever you see it with the value you are supposed to evaluate it at. g(0) = -2x+5 = -2(0)+5 = etc.

OpenStudy (bekkah323):

so g(0)=-2(0)+5 and g(-1)=-2(-1)+5 @whpalmer4

OpenStudy (bekkah323):

now what do i do?

OpenStudy (whpalmer4):

Just simplify your answers